OpenStudy (anonymous):

Kristin opens a savings account with $3500. She deposits $2400 every year into the account that has a 0.65% interest rate, compounded quarterly. If she doesn't withdraw any money, what will the account balance be in 10 years?

OpenStudy (anonymous):

hi i think : fv = 3500 * (1 + 0.0065) ^ 10 + (800*( ( 1 + 0.001625)^ 36 )-1) / 0.001625 fv = 3734 + 29634 = 33,368 $

OpenStudy (anonymous):

thanks!!!